# LifeGPT: Topology-Agnostic Generative Pretrained Transformer Model for Cellular Automata

### v1

@@ 2.3 Training Data Ordering Effects
While the difference in the performance of both models is stark, these findings allow us to draw two significant conclusions. The first is that the fundamental state-transition asymmetry in Life (state-0 and state-1 cells do not follow reciprocal rulesets) impacted the learning process. It appeared that it was ‘easier’ for the model to learn how to predict NGS when many 1-state cells were in the IC, suggesting that the ‘overcrowding’ and ‘stay living’ rules (see the lower-central and rightmost arrows in Fig. 1) was more learnable than the ‘birth’ rule (see the upper-central arrow in Fig. 1). The return to reasonably high accuracy for $\langle\eta\rangle_{\mathrm{IC}}\in\sim[0,0.1)$ suggests that the ‘stay dead’ rule (see the leftmost arrow in Fig. 1) was also reasonably learnable. One possible reason for the model’s inability to extrapolate the birth rule could be that this rule has the most specific conditions for being enacted, requiring a 0-state cell to have exactly 3 1-state neighbors – for a high-entropy training set, the model is simply going to encounter instances of the other 3 rules more frequently. As shown in Fig. 5A, there was still some variation in the measured order parameters of samples in the training set (the distribution was binomial because the data generation process was stochastic in the same way as flipping a fair coin – see section 4.3.5). Still, this variation was clearly not enough to provide the diversity in IC ordering necessary to enable LifeGPT to capture all of the rules of Life accurately.
@@ 2.4 Zero/Few Shot Learning Abilities
Zero-shot and few-shot learning are demonstrated when models perform well on tasks that either do not appear or rarely appear in training data. We show that our training data, containing 10,000 stochastically-generated ICs (see 4.3.7), represents only a small fraction of all possible ICs. It is very unlikely for one or more of the ICs in our testing set (excluding the two samples generated with $p_{x}=0$ and $p_{x}=1$ respectively) to exactly match one or more ICs in our training set. Given that the total number of possible initial conditions is $N_{\mathrm{Total}}=2^{32}$, the probability of randomly testing the model on a training sample is $\frac{10,000}{2^{32}}\approx 2.33\times 10^{-6}=0.000233\%$. We also empirically confirmed that there were no matches between the 8 testing sample ICs (again, excluding $p_{x}=0$ and $p_{x}=1$) and the 10,000 sample training set we used, using a simple search algorithm.
Based on our performance characterization (Fig. 2), LifeGPT displays impressive zero/few-shot learning capabilities. LifeGPT is able to simulate the rules of Life with near-perfect accuracy for a 32 $\times$ 32 toroidal grid by learning off of less than 0.0003% of all possible starting configurations. Still, extremely rare instances of incorrect token outputs are observed, even with a sampling temperature of 0. This is observed experimentally when feeding the output of LifeGPT back into its input, creating two layers of autoregression (we called this method ARAR, for ‘autoregressive autoregressor’ – see section 4.8), which recursively implement the rules of Life. Specifically, an error in a the output of a single cell will cause a perturbation to the Life system, that, over several additional iterations, is highly likely to drastically affect the entire game state with respect to the ground truth. This occured when inputting the ‘Order_param: 0.25’ ($\langle\eta\rangle_{\mathrm{IC}}=0.25$), ‘glider,’ and ‘r-Pentomino’ ICs from the testing into the ARAR algorithm (see $\langle\eta\rangle_{\mathrm{IC}}=0.25$ Animation at Temperature 0 – 250 Timesteps, Glider Animation at Temperature 0 – 250 Timesteps, and r-Pentomino Animation at Temperature 0 – 250 Timesteps, respectively). For the $\langle\eta\rangle_{\mathrm{IC}}=0.25$ and ‘r-Pentomino’ ICs, only a single cell was incorrectly predicted during the generation step which caused divergence from ground truth, while for the ‘glider’ IC, divergence was observed due to an iteration wich incorrectly predicted 2 cells (Fig. 6).
@@ 2.5 Future Work
Future LifeGPTs could be trained on a wide space of CA rulesets, may be able to solve an inverse problem. Here, instead of predicting the precise evolution of CA systems, they may be capable of assigning a representative CA ruleset to a real-world dynamic system, based on data representing the system’s spatiotemporal evolution. This rules-based representation would help advanced the phenomenological understanding of the system at hand, and would enable generative exploration of the system’s configuration-space via IC-variation and subsequent recursive rule application. For instance, this strategy could be applied to better understand the growth-dynamics of plants and fungi, since CA have already been tried as models for many such systems[44, 45, 46, 47, 48], and the growth of such organisms is largely governed by the emergent effects of cellular interactions [49, 50, 51, 52].
@@
Yet another opportunity may arise from the fact that Life has been shown to be Turing complete, via the implementation of a Universal Turing Machine (UTM)[53]. Turing completeness is, in most cases, synonymous with computational universality[54]. If a system can be proven to be Turing complete, then that a system is capable of simulating any Turing machine, or in other words, any classical computer, granted the system is large enough to store the necessary code[55]. Therefore, our work empirically shows that, theoretically, GPT models could become capable of simulating any classical computer algorithm (granted sufficient hardware). This falls in line with theoretical work that has proven the Turing-completeness of the transformer architecture in the past[56, 57]. While our GPT model was trained on game grids far too small to implement a full-scale UTM, and modern-day hardware still limits the practicality of synthesizing a GPT with Life to simulate a UTM, our work nevertheless showcases the possibility for future models to synthesize stochastic generative capabilities with deterministic computational capabilities. This would be particularly useful in the context of hallucination-prone LLM chatbots[58, 59], as simulating CA-based computers might enable them to perform deterministic calculations without compromising their creative, generative capabilities.
@@
Considering that LifeGPT does, occasionally, predict incorrect tokens, more work could be done to investigate other methods of training GPTs on this front. In recent years, more work has been done investigating the potential of transformer models, particularly LLMs, for strategic game-play[23]. In particular, methods which incorporate some kind of reinforcement learning (RL), which involves the LLM making decisions within and getting feedback from a simulated game environment, have been shown to be particularly successful. For these types of applications, transformer models – specifically, LLMs – act as agents which can make a series of decisions, often producing more than one output at a time, conveying several viable options. An RL algorithm works in tandem with the LLM to choose the best decisions from generated options. By playing a given strategy game repeatedly against other AI agents, where each decision leads to a generated reward enabling RL training, the agents will learn over time to take optimal decisions. This strategy usually requires a pretrained LLM chatbot, and does not necessarily involve any LLM fine-tuning steps. Xu et al. (2024) found that the addition of an RL algorithm, as opposed to only using an LLM, helps the model overcome intrinsic biases found in LLMs such as ChatGPT. For a so-called ‘solved game,’ this strategy can ensure that the model’s actions more closely align with the corresponding Nash equilibrium [22].
@@ 3 Conclusion
Future work could also focus on mitigating the effects of occasional errors. The intrinsic stochasticity of the transformer architecture leaves the door open for building better understandings of probabilistic system, and of coarse-grained approximations of the real-world. The topology-agnostic nature of LifeGPT, unlike previous methods of learning Life and other CA, may allow future exploration of a diverse space of nonlinear, dynamical, continuum systems, regardless of the details of spatiotemporal dependencies
@@ 4 Materials and Methods
Codes, data, and relevant animations/figures are available at https://github.com/lamm-mit/LifeGPT
@@ 4.3.6 Validation Set Generation
Validation sets were generated using the same methods in section 4.3.5, as the random.random() function ensures sufficiently random IC generation, ensuring training and validation sets remained entirely independent. Combined with the incredibly large space of possible 32 $\times$ 32 binary arrays ($2962^{32}=4,294,967,$ 296 unique possibilities), this made the likelihood of even a single sample being identical between a 10,000-sample training set and a 1000-sample validation set very low (see section 2.4). This, in turn, ensured that over the course of model training, training loss and validation loss remained independent of each another.
@@ 4.8 Autoregressive Autoregressor Implementation
We also ran the ARAR script (and the Life algorithm) for 249 iterations (resulting in 250 game states, including the ICs), using only the epoch 50, temperature=0 version of LifeGPT, due to time and compute constraints, for all 10 samples in the testing set. For each game state, we compared LifeGPT’s predictions to the GT Life algorithm’s output using the metric ‘Error Rate’, defined as:
@@ Acknowledgments
Figure 6: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ’Glider’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 7: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ’Cloverleaf’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 8: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ’Hammerhead Spaceship’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 9: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ’Blinkers’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 10: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ’r-Pentomino’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.

### v2

@@ 2.3 Training Data Ordering Effects
While the difference in the performance of both models is stark, these findings allow us to draw two significant conclusions. The first is that the fundamental state-transition asymmetry in Life (state-0 and state-1 cells do not follow reciprocal rulesets) impacted the learning process. It appeared that it was ‘easier’ for the model to learn how to predict NGS when many 1-state cells were in the IC, suggesting that the ‘overcrowding’ and ‘stay living’ rules (see the lower-central and rightmost arrows in Fig. 1) was more learnable than the ‘birth’ rule (see the upper-central arrow in Fig. 1). The return to reasonably high accuracy for $\langle\eta\rangle_{\mathrm{IC}}\in\sim[0,0.1)$ suggests that the ‘stay dead’ rule (see the leftmost arrow in Fig. 1) was also reasonably learnable. One possible reason for the model’s inability to extrapolate the birth rule could be that this rule has the most specific conditions for being enacted, requiring a 0-state cell to have exactly 3 1-state neighbors – for a high-entropy training set, the model is simply going to encounter instances of the other 3 rules more frequently. As shown in Fig. 5A, there was still some variation in the measured order parameters of samples in the training set (the distribution was binomial because the data generation process was stochastic in the same way as flipping a fair coin – see section 4.3.5). Still, this variation was clearly not enough to provide the diversity in IC ordering necessary to enable LifeGPT to capture all of the rules of Life accurately. This version of LifeGPT effectively found a ‘pocket of computational reducibility’ [8] within Life which applied only to ICs outside of a particular order parameter range.
@@ 2.4 Zero/Few Shot Learning Abilities
Zero-shot and few-shot learning are demonstrated when models perform well on tasks that either do not appear or rarely appear in training data. We show that our training data, containing 10,000 stochastically-generated ICs (see 4.3.7), represents only a small fraction of all possible ICs. It is extremely improbable for one or more of the ICs in our testing set (excluding the two samples generated with $p_{x}=0$ and $p_{x}=1$ respectively) to exactly match one or more ICs in our training set. Given that the total number of possible initial conditions is $N_{\mathrm{Total}}=2^{1024}\approx 1.80\times 10^{308}$, the probability of randomly testing the model on a training sample is $\frac{10,000}{2^{1024}}\approx 5.56\times 10^{-305}$. We also empirically confirmed that there were no matches between the 8 testing sample ICs (again, excluding $p_{x}=0$ and $p_{x}=1$) and the 10,000 sample training set we used, using a simple search algorithm.
Based on our performance characterization (Fig. 2), LifeGPT displays impressive zero/few-shot learning capabilities. LifeGPT is able to simulate the rules of Life with near-perfect accuracy for a 32 $\times$ 32 toroidal grid by learning off of less than one part in one cubic-googol (i.e. $<10^{-300}$) of all possible starting configurations. Still, extremely rare instances of incorrect token outputs are observed, even with a sampling temperature of 0. This is observed experimentally when feeding the output of LifeGPT back into its input, creating two layers of autoregression (we called this method ARAR, for ‘autoregressive autoregressor’ – see section 4.8), which recursively implement the rules of Life. Specifically, an error in a the output of a single cell will cause a perturbation to the Life system, that, over several additional iterations, is highly likely to drastically affect the entire game state with respect to the ground truth. This occurred when inputting the ‘Order_param: 0.25’ ($\langle\eta\rangle_{\mathrm{IC}}=0.25$), ‘glider,’ and ‘r-Pentomino’ ICs from the testing into the ARAR algorithm (see $\langle\eta\rangle_{\mathrm{IC}}=0.25$ Animation at Temperature 0 – 250 Timesteps, Glider Animation at Temperature 0 – 250 Timesteps, and r-Pentomino Animation at Temperature 0 – 250 Timesteps, respectively). For the $\langle\eta\rangle_{\mathrm{IC}}=0.25$ and ‘r-Pentomino’ ICs, only a single cell was incorrectly predicted during the generation step which caused divergence from ground truth, while for the ‘glider’ IC, divergence was observed due to an iteration wich incorrectly predicted 2 cells (Fig. 6).
@@ 2.5 Future Work
Future LifeGPT-like models, trained on a massive space of CA rulesets (akin to a tiny piece of the ‘ruliad’ [48]), may be able to solve a multitude of inverse problems. Here, instead of predicting the precise evolution of CA systems, they may be capable of assigning a representative CA ruleset to a real-world dynamical system to raw data representing the system’s spatiotemporal evolution. This rules-based representation would help advanced the phenomenological understanding of the system at hand, as different coarse-grained rules could be uncovered for different length-scales, and would enable generative exploration of the system’s configuration-space via IC-variation and subsequent recursive rule application. For instance, this strategy could be applied to better understand the growth-dynamics of plants and fungi, since CA have already been tried as models for many such systems [49, 50, 51, 52, 53], and the growth of such organisms is largely governed by the emergent effects of cellular interactions [54, 55, 56, 57]. This concept echoes the existing idea that AIs might be able to find “pockets of computational reducibility” in otherwise computationally irreducible systems.Some theorize that transformer models could prune search spaces in multicomputational systems, where non-deterministic state transitions explore multiple paths to reach final states. This raises the question of whether AIs could similarly find rule sets that meet desired emergent behaviors, with the answer depending on the structure of the rule space. One hypothesis on this subject suggests that higher-dimensional rule spaces may allow enough emergent complexity for random rule changes to effectively escape local minima on cost surfaces [8].
@@
Yet another opportunity may arise from the fact that Life has been shown to be Turing complete, via the implementation of a Universal Turing Machine (UTM) [58]. Turing completeness is, in most cases, synonymous with computational universality[59]. If a system can be proven to be Turing complete, then that a system is capable of simulating any Turing machine, or in other words, any classical computer, granted the system is large enough to store the necessary code [60]. Therefore, our work empirically shows that, theoretically, GPT models could become capable of simulating any classical computer algorithm (granted sufficient hardware). This falls in line with theoretical work that has proven the Turing-completeness/computational universality of the transformer architecture  [61, 62] as well as autoregressive LLMs specifically [63]. While our GPT model was trained on game grids far too small to implement a full-scale UTM within Life, and modern-day hardware still limits the practicality of synthesizing a GPT with Life to simulate a UTM, our work nevertheless showcases the possibility for future models to synthesize stochastic generative capabilities with deterministic computational capabilities. This would be particularly useful in the context of hallucination-prone LLM chatbots [64, 65], as simulating CA-based computers might enable them to perform deterministic calculations without compromising their creative, generative capabilities.
@@
Considering that LifeGPT does, occasionally, predict incorrect tokens, more work could be done to investigate other methods of training GPTs on this front. In recent years, researchers have investigated the potential of transformer models – particularly LLMs – to excel at strategic game-play [26]. In particular, methods which incorporate some kind of reinforcement learning (RL), which involves the LLM making decisions within and getting feedback from a simulated game environment, have been shown to be particularly successful. For these types of applications, transformer models – specifically, LLMs – act as agents which can make a series of decisions, often producing more than one output at a time, conveying several viable options. An RL algorithm works in tandem with the LLM to choose the best decisions from generated options. By playing a given strategy game repeatedly against other AI agents, where each decision leads to a generated reward enabling RL training, the agents will learn over time to take optimal decisions. This strategy usually requires a pretrained LLM chatbot, and does not necessarily involve any LLM fine-tuning steps. Xu et al. (2024) found that the addition of an RL algorithm, as opposed to only using an LLM, helps the model overcome intrinsic biases found in LLMs such as ChatGPT. For a so-called ‘solved game,’ this strategy can ensure that the model’s actions more closely align with the corresponding Nash equilibrium [25].
@@ 3 Conclusion
Future work could also focus on mitigating the effects of occasional errors. The intrinsic stochasticity of the transformer architecture leaves the door open for building better understandings of probabilistic systems, and of coarse-grained approximations of the real-world. The topology-agnostic nature of LifeGPT, unlike previous methods of learning Life and other CA, may allow future exploration of a diverse space of nonlinear, dynamical, continuum systems, regardless of the details of spatiotemporal dependencies.
@@ 4 Materials and Methods
Codes, data, and relevant animations/figures are available at https://github.com/lamm-mit/LifeGPT.
@@ 4.3.6 Validation Set Generation
Validation sets were generated using the same methods in section 4.3.5, as the random.random() function ensures sufficiently random IC generation, ensuring training and validation sets remained entirely independent. Combined with the incredibly large space of possible 32 $\times$ 32 binary arrays ($2^{32\times 32}\approx 1.80\times 10^{308}$ unique possibilities), this made the likelihood of even a single sample being identical between a 10,000-sample training set and a 1000-sample validation set negligible (see section 2.4). This, in turn, ensured that over the course of model training, training loss and validation loss remained independent of one another.
@@ 4.8 Autoregressive Autoregressor Implementation
We also ran the ARAR script (and the Life algorithm) for 249 iterations (resulting in 250 game states, including the ICs), using only the epoch 50, temperature=0 version of LifeGPT due to time and compute constraints, for all 10 samples in the testing set. For each game state, we compared LifeGPT’s predictions to the GT Life algorithm’s output using the metric ‘Error Rate’, defined as:
@@ Acknowledgments
Figure 6: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ‘Glider’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 7: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ‘Cloverleaf’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 8: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ‘Hammerhead Spaceship’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 9: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ‘Blinkers’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.
Figure 10: Top row: the ground truth (GT) progression of Life, for an IC corresponding to the ‘r-Pentomino’ pattern. Middle row: LifeGPT’s best prediction (Pred) using the ARAR method at temperature=0. Bottom row: the discrepancy between the ground truth and LifeGPT’s predictions (Error), where blue indicates no discrepancy, and yellow indicates an incorrect cell predicted by LifeGPT.